# Scaling Limit of a Stochastic Clustering Model on $\mathds{R}$

### v1

@@ 4 A Further Conjecture
The conjectured convergence of the measures $\overleftarrow{F}^{(t)}$ should follow from the pointwise convergence of their Laplace transforms. Indeed, define for $s\geqslant 0$,
$G^{(t)}(s):=\int_{\mathds{R}}e^{-sx}dF^{(t)}(x).$
Using the dynamics for $F$, we get:
$G^{(t+1)}(s)$ $=\int_{\mathds{R}}e^{-sx}dF^{(t+1)}(x)$
$=\int_{\mathds{R}}e^{-s(x+\beta^{(t+1)}_{x})}dF^{(t+1)}(x+\beta^{(t+1)}_{x})$
$=\int_{\mathds{R}}e^{-sx-s\beta^{(t+1)}_{x}}dF^{(t)}(x)+\int_{\mathds{R}}e^{-sx-s\beta^{(t+1)}_{x}}d\alpha^{(t+1)}_{x}$
$=:H_{\beta}^{(t+1)}(s)+H_{\alpha}^{(t+1)}(s),$
where we treat $\alpha^{(t)}(x)$ as a signed measure. Here, the second equality follows since the mapping $x\mapsto x+\beta^{(t+1)}_{x}$ is monotone increasing in $x$: we have $x+\beta^{(t+1)}_{x}=\frac{2}{3}x+N_{\rho}^{(t+1)}(x)$, where $N_{\rho}^{(t+1)}(\cdot)$ accounts for any positive jumps due to $\rho^{(t+1)}$.
For each (fixed) $s$, we consider the stochastic process $G^{(t)}(s)$. To do this, one could analyze the two functions $H_{\beta}^{(t+1)}(s),\ H_{\alpha}^{(t+1)}(s)$ separately as follows. Notice that $\beta^{(t)}_{x}$ and $\alpha^{(t)}_{x}$ are (conditionally) mean-zero for each time $t$ and location $x$. The simpler term to analyze is $H_{\beta}^{(t+1)}(s)$. We have
$H_{\beta}^{(t+1)}(s)$ $=\int_{\mathds{R}}e^{-sx-s\beta^{(t+1)}_{x}}dF^{(t)}(x)=\int_{\mathds{R}}e^{-s\beta^{(t+1)}_{x}}e^{-sx}dF^{(t)}(x),$
and with a suitable MGF control, we expect that
$H_{\beta}^{(t+1)}(s)\leqslant C^{(t)}\cdot G^{(t)}(s),$
for some constants satisfying $\prod_{i=1}^{\infty}C^{(t)}<\infty$. Here, we use the fact that $\rho^{(t+1)}(x)$ is independent of the canonical filtration (up to time $t$). The term $H_{\alpha}^{(t+1)}(s)$ is more difficult to analyze due to the fact that the $\alpha^{(t+1)}_{\cdot}$ process is highly correlated with the $\beta^{(t+1)}_{\cdot}$ process and will be analyzed in a future paper.
4.1 Simulations of Conjecture 1
In Figure 5, we show simulations for two independent copies of our dynamics. We observe the conjectured convergence as $t$ increases. However, we obviously cannot pictorially depict the convergence of the measure over distribution functions.
(a) First Simulation
(b) Second Simulation
@@ 5 Future Work
In this paper, we analyze a simple dynamical clustering algorithm for $\mathds{R}$-valued data in which each data point attempts to cluster with either its left or right neighbor, chosen uniformly at random. For this model, we establish (under mild assumptions) a unique scaling limit which is independent of the initial data.
We conjecture that (under similar assumptions), a similar scaling limit holds for other realistic dynamic clustering algorithms, such as Algorithm 2 from the introduction of this paper. For this algorithm, the scaling limit appears to depend on the initial data. The proof technique used in this work does not apply to Algorithm 2, as the weight sequences in the time reversal are no longer described via integer weights. Determining the existence of stationary distributions for this type of dynamics remains an important area of future work. New techniques are required to study these algorithms in a greater generality.

### v2

@@ 5 Proofs of Technical Results
5.2 Proof of Lemma 3.7
We will use Lemma 3.8. Let $Y\sim\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$ be the renewal gap distribution with pmf $f_{k}=k\cdot 2^{-k-1}\text{ for }k\in\mathds{N},$ mean $\mu=3$, and mgf
$\varphi(\theta)=\operatorname{\mathds{E}}e^{\theta Y}=e^{\theta}(2-e^{\theta})^{-2}\geqslant e^{\mu\theta}\text{ for }\theta<\ln 2.$
Let $Y_{*}$ denote a random variable with the equilibrium residual lifetime distribution with pmf
$f^{\star}_{k}=\frac{1}{\mu}\operatorname{\mathds{P}}(Y\geqslant k)=\frac{1}{3}(k+1)\cdot 2^{-k}\text{ for }k\geqslant 1.$
It is easy to check that both $Y$ and $Y^{\star}$ satisfy the condition in Lemma 3.9. Thus, we have $C_{\alpha}=\mathds{1}_{\alpha\geqslant 0}+e^{\theta-\alpha}\mathds{1}_{\alpha<0}$ when $\alpha=\varphi^{-1}(e^{\theta})\leqslant\theta/\mu$. Similarly, $C^{\star}_{\alpha}=\mathds{1}_{\alpha\geqslant 0}+\operatorname{\mathds{E}}e^{\alpha(Y^{\star}-1)}\mathds{1}_{\alpha<0}$.
Next, we solve for $\varphi(x)=e^{\theta}$, to get
$e^{2x}-(4+e^{-\theta})e^{x}+4=0\text{ or }e^{x}=\frac{4+e^{-\theta}-\sqrt{e^{-\theta}(e^{-\theta}+8)}}{2}.$
Thus, for $\theta\in\mathds{R}$, we have
$\psi(\theta)$ $=\theta/3-\varphi^{-1}(e^{\theta})$
$=\theta/3-\log\left(1-\frac{2(e^{-\theta}-1)}{2+e^{-\theta}+\sqrt{e^{-\theta}(e^{-\theta}+8)}}\right)$
$\leqslant\theta/3-\log\left(1-\frac{e^{-\theta}-1}{2+e^{-\theta}}\right)\leqslant\log\left(\frac{2}{3}e^{\theta/3}+\frac{1}{3}e^{-2\theta/3}\right)\leqslant\frac{1}{8}\theta^{2}.$
Using Lemma 3.8, we get
$\operatorname{\mathds{E}}e^{\theta(x/\mu-N_{\rho}^{\star}(x))}$ $\leqslant(1+A\left|\theta\right|)\cdot e^{x\theta^{2}/8}$
$\text{ and }\operatorname{\mathds{E}}e^{\theta(x/\mu-N_{\rho}^{\star}(x))}\mathds{1}_{{x\in\rho}}$ $\leqslant e^{\max(0,-\theta)}\cdot\frac{1}{\mu}e^{x\theta^{2}/8}$
$\text{ for all }x\geqslant 1,\theta\in\mathds{R}.$ This proves the lemma. $\blacksquare$
5.3 Proof of Lemma 3.8
Fix $\theta\in\mathds{R}$, and define $m_{t}:=\operatorname{\mathds{E}}e^{-\theta N_{t}},m^{\star}_{t}:=\operatorname{\mathds{E}}e^{-\theta N^{\star}_{t}}\text{ for }t\geqslant 0.$ We have
$m_{t}=\operatorname{\mathds{P}}(Y>t)+e^{-\theta}\sum_{k=1}^{t}f_{k}m_{t-k}.$
Define
$\alpha:=\varphi^{-1}(e^{\theta})\leqslant\theta/\mu$
so that $\varphi(\alpha)=e^{\theta}$. The last inequality follows from Jensen’s inequality. Note that $\alpha>0$ for $\theta>0$ and $\alpha<0$ for $\theta<0$. Define $\overline{m}_{t}:=e^{\alpha t}m_{t}.$ Let $f^{(\alpha)}$ be the pmf given by $f^{(\alpha)}_{k}:=e^{\alpha k-\theta}f_{k},k\geqslant 1$ and $Y^{(\alpha)}$ be a random variable with pmf $f^{(\alpha)}$. We get
$\overline{m}_{t}=\mu e^{\alpha t}f_{t+1}^{\star}+\sum_{k=1}^{t}f^{(\alpha)}_{k}\cdot\overline{m}_{t-k}.$
Let
$A(\alpha):=\sum_{t=0}^{\infty}e^{\alpha t}f_{t+1}^{\star}=e^{-\alpha}\operatorname{\mathds{E}}e^{{\alpha Y^{\star}}}$
and $Z_{\alpha}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{*,(\alpha)}-1$ be a random variable with pmf $e^{\alpha t}f_{t+1}^{\star}/A(\alpha)$, $t\geqslant 0$. Thus, we have
$\overline{m}_{t}$ $=\mu A(\alpha)\cdot\operatorname{\mathds{P}}(Y^{*,(\alpha)}+Y_{1}^{(\alpha)}+Y_{2}^{(\alpha)}+\cdots+Y_{k}^{(\alpha)}=t+1\text{ for some }k\geqslant 0)$
$\leqslant\frac{1}{C_{\alpha}}\cdot e^{\theta-\alpha}\cdot\operatorname{\mathds{E}}Y^{(\alpha)}\cdot\operatorname{\mathds{P}}(Y^{(\alpha),*}+Y_{1}^{(\alpha)}+Y_{2}^{(\alpha)}+\cdots+Y_{k}^{(\alpha)}=t+1\text{ for some }k\geqslant 0)$
$\leqslant\frac{1}{C_{\alpha}}\cdot e^{\theta-\alpha}.$
Here, we used the fact that for all $k\geqslant 1$,
$\frac{\mu A(\alpha)\cdot\operatorname{\mathds{P}}(Y^{*,(\alpha)}=k)}{\operatorname{\mathds{E}}Y^{(\alpha)}\cdot\operatorname{\mathds{P}}(Y^{(\alpha),*}=k)}=\frac{e^{\theta-\alpha}\operatorname{\mathds{P}}(Y\geqslant k)}{\sum_{t\geqslant k}e^{\alpha(t-k)}f_{t}}\leqslant e^{\theta-\alpha}\cdot\frac{1}{C_{\alpha}}.$
Similarly, for the equilibrium renewal process, we get $m^{\star}_{t}=\operatorname{\mathds{P}}(Y^{\star}>t)+e^{-\theta}\sum_{k=1}^{t}f^{\star}_{k}m_{t-k}$, or
$e^{\alpha t}m^{\star}_{t}$ $=e^{\alpha t}\operatorname{\mathds{P}}(Y^{\star}>t)+e^{-\theta}\sum_{k=1}^{t}e^{\alpha k}f^{\star}_{k}\cdot\overline{m}_{t-k}$
$\leqslant e^{-\alpha}\cdot e^{\alpha(t+1)}\operatorname{\mathds{P}}(Y^{\star}\geqslant t+1)+\frac{1}{C_{\alpha}}\cdot e^{-\alpha}\sum_{k=1}^{t}e^{\alpha k}f^{\star}_{k}$
$\leqslant\max\left(\frac{1}{C^{\star}_{\alpha}},\frac{1}{C_{\alpha}}\right)\cdot\operatorname{\mathds{E}}e^{\alpha(Y^{\star}-1)}.$
For the second part, define
$n_{t}:=\mu e^{\alpha t}\operatorname{\mathds{E}}(e^{\theta N^{\star}_{t}}\cdot\mathds{1}_{t\in\rho^{\star}}),t\geqslant 1,$
which satisfies the recursion
$n_{t}$ $=\mu e^{-\theta}e^{\alpha t}f_{t}^{\star}+\sum_{k=1}^{t}f^{(\alpha)}_{k}\cdot n_{t-k}.$
This is bounded by $1/C_{\alpha}$ by the same argument used in the first part. Simplifying, we get the result. $\blacksquare$
5.4 Proof of Lemma 3.9
The first equality follows from the fact that $\operatorname{\mathds{P}}(Y^{\star}=k)=\frac{1}{\mu}\operatorname{\mathds{P}}(Y\geqslant k),k\geqslant 1$ and by changing the order of the sums. For the second result, we will show that $\operatorname{\mathds{E}}(e^{\alpha(Y-k)}\mid Y\geqslant k)$ is increasing in $k\geqslant 1$, when $\alpha<0$. Then the result follows by taking $k=1$. Finally, we have
$e^{\alpha(k+1)}\operatorname{\mathds{P}}(Y\geqslant k)\operatorname{\mathds{P}}(Y\geqslant k+1)\cdot$ $\left(\operatorname{\mathds{E}}(e^{\alpha(Y-k-1)}\mid Y\geqslant k+1)-\operatorname{\mathds{E}}(e^{\alpha(Y-k)}\mid Y\geqslant k)\right)$ (5.1)
$=\sum_{i=k}^{\infty}f_{i}\cdot\sum_{j=k+1}^{\infty}e^{\alpha j}f_{j}-e^{\alpha}\sum_{i=k+1}^{\infty}f_{i}\cdot\sum_{j=k}^{\infty}e^{\alpha j}f_{j}.$
Simplifying, we get
(5.1) $=(1-e^{\alpha})\cdot\sum_{i=k}^{\infty}f_{i}\cdot\sum_{j=k}^{\infty}e^{\alpha j}f_{j}-f_{k}\sum_{i=k}^{\infty}f_{i}(1-e^{\alpha})\sum_{j=k}^{i}e^{\alpha j}$
$=(1-e^{\alpha})\cdot\sum_{j=k}^{\infty}e^{\alpha j}\left(f_{j}\operatorname{\mathds{P}}(Y\geqslant k)-f_{k}\operatorname{\mathds{P}}(Y\geqslant j)\right)>0.$
This completes the proof. $\blacksquare$
5.5 Proof of Lemma 3.10
Define $\xi_{k}:=\mathds{1}_{k\in\rho^{\star}}-1/\mu,$ so that $(\xi_{k})_{k\in\mathds{Z}}$ is stationary, centered, and bounded by $|\xi_{k}|\leqslant 1$. Recall that $Y_{*}$, the equilibrium residual lifetime distribution, satisfies $\operatorname{\mathds{P}}(Y_{*}\geqslant k)\leqslant c_{0}e^{-\theta k}$, $k\geqslant 1$, for some $\theta>0$. For completeness, we present a proof sketch based on exponential mixing of the random variables $(\xi_{k})$ and bounds on the joint cumulants. Note that all of the tools used are available in the literature. One can also derive a similar result using block decomposition.
Using the regenerative property of the renewal process, the occurrence of a renewal in a separating interval cuts the dependence between the past and the future. It follows that $(\xi_{k})$ is exponentially $\beta$-mixing, i.e.,
$\beta(r)\leqslant C_{1}e^{-c_{1}r},\quad r\geqslant 1,$
for some constants $C_{1},c_{1}>0$ depending only on the law of $Y$ (see, e.g.,, the mixing background in [8] and regeneration coupling arguments in [31]).
Next, standard cumulant estimates for bounded exponentially mixing sequences (cf. [9, 17]) imply that there exist constants $A,B>0$ such that, for every $m\geqslant 2$ and every $i_{1},\ldots,i_{m}\in\mathds{Z}$,
$\big|\kappa(\xi_{i_{1}},\ldots,\xi_{i_{m}})\big|\leqslant A^{m-1}(m-1)!\,e^{-B\,T(i_{1},\ldots,i_{m})},$
where $\kappa(\cdot)$ denotes joint cumulant and $T(i_{1},\dots,i_{m})$ is the minimal spanning-tree length of $\{i_{1},\ldots,i_{m}\}$ in the graph metric on $\mathds{Z}$. By the multi-linearity of cumulants,
$\kappa_{m}(X_{t})=\sum_{i_{1},\dots,i_{m}=1}^{t}v_{i_{1}}\cdots v_{i_{m}}\,\kappa(\xi_{i_{1}},\ldots,\xi_{i_{m}}),\qquad m\geqslant 1,$
and $\kappa_{1}(X_{t})=\operatorname{\mathds{E}}X_{t}=0$. Combining the preceding bound with the exponential summability of the kernel $e^{-B|i-j|}$ yields
$|\kappa_{m}(X_{t})|\leqslant K^{m-1}m!\left\|\boldsymbol{v}\right\|_{2}^{2}\left\|\boldsymbol{v}\right\|_{\infty}^{m-2},\qquad m\geqslant 2,$
for $K=4A\sum_{r=0}^{\infty}e^{-Br}$. Finally, for $|\lambda|\left\|\boldsymbol{v}\right\|_{\infty}<(2K)^{-1}$, the cumulant expansion of the log-moment generating function converges absolutely
$\log\operatorname{\mathds{E}}e^{\lambda X_{t}}=\sum_{m\geqslant 2}\frac{\kappa_{m}(X_{t})}{m!}\lambda^{m}.$
Hence
$\log\operatorname{\mathds{E}}e^{\lambda X_{t}}\leqslant K\left\|\boldsymbol{v}\right\|_{2}^{2}\sum_{m\geqslant 2}(K|\lambda|\left\|\boldsymbol{v}\right\|_{\infty})^{m-2}\lambda^{2}=2K\lambda^{2}\left\|\boldsymbol{v}\right\|_{2}^{2}.$
Renaming constants gives the claimed MGF bound. $\blacksquare$
@@ 6 Future Work
In this paper, we analyze a simple dynamical clustering algorithm for $\mathds{R}$-valued data in which each data point attempts to cluster with either its left or right neighbor, chosen uniformly at random. For this model, we establish (under mild assumptions) a unique scaling limit that is independent of the initial data.
We conjecture that (under similar assumptions), a similar scaling limit holds for other realistic dynamic clustering algorithms, such as Algorithm 2 from the introduction of this paper. For this algorithm, the scaling limit appears to depend on the initial data. The proof technique used in this work does not apply to Algorithm 2, since the weight sequences under time reversal are no longer integer-valued. Determining the existence of stationary distributions for this type of dynamics remains an important area of future work. New techniques are required to study these algorithms with greater generality.
Also, an open problem is to identify the limiting non-renewal point process as the unique fixed point of an appropriate stochastic dynamics. Based on simulations, it seems that the weak limit (in the local weak sense) of the genealogy tree exists. Another relevant question is whether we can embed it in a hyperbolic space and obtain a space-time limit that includes both the spatial component and the time dynamics.